HETAV PATEL

Grade 11,

The precession frequency of a proton in a magnetic field of 0.5Weber/m 2

The precession frequency of a proton in a magnetic field of 0.5Weber/m2

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Dear student 
The precessional frequency of nuclei of a substance placed in a static magnetic field B0 is calculated from the Larmor Equation:
ω = γB
The proton's gyromagnetic ratio, symbol γp, is 2.675222005×10^8 rad⋅s−1⋅T−1
ω = 2.675222005×10^8 * 0.5 = 1.3375 *10^8 
Hope this helps